The U.S. Energy Department states that 60% of all U.S. households have ceiling fans. In addition, 29% of all U.S. households have an outdoor grill. Suppose 13% of all U.S. households have both a ceiling fan and an outdoor grill. A U.S. household is randomly selected.

a. What is the probability that the household has a ceiling fan or an outdoor grill?

b. What is the probability that the household has neither a ceiling fan nor an outdoor grill?

c. What is the probability that the household does not have a ceiling fan and does have an outdoor grill?

d. What is the probability that the household does have a ceiling fan and does not have an outdoor grill?



1
Expert's answer
2021-06-09T15:56:01-0400

Let "C"  represent that US household has ceiling fan: "P(C)=0.60."

Let "G"  represent that US household has outdoor grill: "P(G)=0.29."

The probability that both ceiling fan and outdoor grill is "P(C\cap G)=0.13."


a.

"P(C\cup G)=P(C)+P(G)-P(C\cap G)"

"=0.60+0.29-0.13=0.76"

The probability that the household has a ceiling fan or an outdoor grill is 0.76.


b.


"P(C'\cap G')=1-P(C\cup G)=1-0.76"




"=0.24"

The probability that the household has neither a ceiling fan nor an outdoor grill IS 0.24.


c.


"P(C'\cap G)=P(G)-P(C\cap G)=0.29-0.13"




"=0.16"

The probability that the household does not have a ceiling fan and does have an outdoor grill is 0.16.


d.


"P(C\cap G')=P(C)-P(C\cap G)=0.60-0.13"

"=0.47"

The probability that the household does have a ceiling fan and does not have an outdoor grill iis 0.47.